OpenStudy (anonymous):

what about tan squared (x+(Pi/4))=3?

OpenStudy (anonymous):

\[\tan^2(x+\pi/4) = 3\] \[\tan (x+\pi/4) = \sqrt{3}\] \[\frac{ 1+\tan x }{ 1-\tan x }= \sqrt{3}\]

OpenStudy (anonymous):

simplify for tanx